OpenStudy (anonymous):

If the graph of 3x-5y=s passes through points (3,1) and (5,r) what is the value of r?? a)2 b) 11/5 c) 16/5 d) 4 e) 25/4 how did you do this?

OpenStudy (anonymous):

1) Substitute (3,1) into the equation of the graph to get a value of s. 2) Substitute the value of s into the 3x-5y=s equation to get a general form of the equation. 3) Substitute (5,r) into the equation to find for r.

OpenStudy (anonymous):

ohhhh so b? :D

OpenStudy (anonymous):

THANK YOU THANK YOU SOSOSO MUCH

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

can u help me with another?

OpenStudy (anonymous):

if g(x)= f(x)+k and the graph of g(x) (greater than or equal to ) 0 for all vlaues of x, which of the following CANNOT be the value of k? 3 4 5 6 7

OpenStudy (anonymous):

those are the choices

OpenStudy (anonymous):

hellooooo???? :(

OpenStudy (anonymous):

oh nvm LOL i see you typingg

OpenStudy (anonymous):

when g(x) is >= 0, means that f(x) + k is >= 0 also. Will you have expressions of fx and gx?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is the answer 7???

OpenStudy (anonymous):

herro???